Question Part
Points
1
–/10
Total
–/10
...Given that the random variable X is normally distributed with a mean of 80 and a standard deviation of 10, P(85 ≤ X ≤ 90) is:
a. 0.5328
b. 0.1498
c. 0.5841
d. 0.3413
e. 0.1915


Answer by stanbon(75887)

Given that the random variable X is normally distributed with a mean of 80 and a standard deviation of 10, P(85 ≤ X ≤ 90) is:
a. 0.5328
b. 0.1498
c. 0.5841
d. 0.3413
e. 0.1915
-----------
z(85) = (85-80)/10 = 5/10 = 1/2
z(90) = (90-80)/10 = 10/10 = 1
----
P(85<= x <=90) = P(1/2<= z <=1) = normalcdf(1/2,1) = 0.1499
====================
Cheers,
Stan H.
====================